OpenStudy (anonymous):

solve for x: x^(1/2) - 5x^(1/4) + 6 = 0

OpenStudy (agreene):

x= 16 and 81

OpenStudy (anonymous):

Can you tell me how you did that?

OpenStudy (agreene):

I could, but you wouldnt like it.

OpenStudy (anonymous):

My bad you are right

OpenStudy (anonymous):

I need to know how you came about the solution, as I don't understand how to do the problem :/

OpenStudy (anonymous):

You can use substitution a=x^(1/4) and a^2=x^(1/2)

OpenStudy (agreene):

Let u = x^(1/4) u^2-5 u+6 = 0 u^2-5 u = -6 u^2-5 u+25/4 = 1/4 (u-5/2)^2 = 1/4 abs(u-5/2) = 1/2 u-5/2 = -1/2 or u-5/2 = 1/2 u = 2 or u-5/2 = 1/2 Replace u = x^(1/4): x^(1/4) = 2 or u-5/2 = 1/2 Let x = r (cos(theta)+i sin(theta)) and write phasor form: (r (cos(theta)+i sin(theta)))^(1/4) = 2 (cos(0)+i sin(0)) Use de Moivre's theorem: r^(1/4) (cos(theta/4)+i sin(theta/4)) = 2 (cos(0)+i sin(0)) Equating moduli and arguments gives: r^(1/4) = 2 and theta/4 = 2 pi k Solutions: r = 16 and theta_1 = 0 Then the equation x^(1/4) = 2 has the solutions: x = 16 or u-5/2 = 1/2 x = 16 or u = 3 Replace u = x^(1/4): x = 16 or x^(1/4) = 3 Let x = r (cos(theta)+i sin(theta)) and write 3 in phasor form: (r (cos(theta)+i sin(theta)))^(1/4) = 3 (cos(0)+i sin(0)) Use de Moivre's theorem: r^(1/4) (cos(theta/4)+i sin(theta/4)) = 3 (cos(0)+i sin(0)) Equating moduli and arguments gives: r^(1/4) = 3 and theta/4 = 2 pi k Solutions: r = 81 and theta_1 = 0 x = 16 or x = 81

OpenStudy (agreene):

Thats how I did it... like I said, you probably wouldnt like it, but it's the quickest method (that i see right now)

OpenStudy (anonymous):

Yeah I see how that works. I just wonder if there is a different way to do it as this is for an Intermediate Algebra class and cosines and sines aren't used at all.

OpenStudy (agreene):

you can try and combine the fractional powers and then evaluate only on the positives and then write a proof that would allow you to operate it on the negative side too, but that seems like a lot more work to me O_o

OpenStudy (anonymous):

Hmm this just doesn't seem to fit in with what is being taught in my class at all.

OpenStudy (agreene):

Well, most algebra classes/books/teachers are stupid and don't realize that : \[\sqrt {x^2} \ne x\] So they do things that I cant, because I know that's the case O_o

OpenStudy (anonymous):

rather than all that sine cosine crap I just put u=3 and u=2 then replaced u back with x^(1/4) then I just did 3^4 and 2^4 to get the same answers

OpenStudy (anonymous):

@Docswoops don't worry. agreene just try to impress you with all his knowledge :-))) You already had an answer.

OpenStudy (anonymous):

In any case I understand how to come to the answer now lol.

OpenStudy (agreene):

@MarinaDL he is trying to find an answer he understands how to get... and I'm trying to think of a solution for him. Don't presume you know what I'm doing at any given time. @sclowers You're probably on the right tract but your Lemma has the incorrect powers :X

OpenStudy (anonymous):

Calm down guys it all makes sense now haha! Honestly I just didn't notice replacing x^(1/4) with "u".... was simple after that.

OpenStudy (agreene):

Thats what I did, but I had to use de Moivres because of the above situation with roots not cancelling for the negative numbers. (else you are needing a proof it hold for the negatives)

OpenStudy (anonymous):

@agreene Thanks for detailed explanation.

OpenStudy (agreene):

But, alas, if you understand it, then meh, more power to yah! lol... I'm probably being a douche right now.

OpenStudy (anonymous):

I'm assuming that level of answer is above my class level.